OpenStudy (gigi1103):

The Mitchells are renting a boat for the day. It costs $100, plus $20 for each hour. They have to pay for a whole hour even if they are not out there for a whole hour. For example if they boat for 3 and a half hours, they have to pay for 4 hours. They don't want to spend more than $250 for the day. How many hours can they boat? Write an inequality and solve. A) 100 + 20x ≤ 250. They can boat for 7 hours. B) 100 + 20x ≥ 250. They can boat for 8 hours. C) 100 + 20x ≤ 250. They can boat for 8 hours. D) 100 + 20x ≥ 250. They can boat for 7 hours.

OpenStudy (gigi1103):

@AlexandervonHumboldt2

OpenStudy (gigi1103):

@crianca

OpenStudy (gigi1103):

@texaschic101

OpenStudy (gigi1103):

@pooja195

OpenStudy (dangerousjesse):

\[250-100+20x=10\]\[250-100=150\]\[150-20x=10\]\[150-10=140\]\[140\div20=x\]

OpenStudy (dangerousjesse):

What is 140/20?

OpenStudy (gigi1103):

140?20 = 7?

OpenStudy (dangerousjesse):

Please don't give out direct answers. We're here to help, not do their work for them.

OpenStudy (texaschic101):

100 + 20x <= 250 20x <= 250 - 100 20x < = 150 x < = 150/20 x <= 7.5...so they have to pay for 8 hrs are you sure it is not C

OpenStudy (dangerousjesse):

Yep, 7 hours :)

OpenStudy (dangerousjesse):

They have a budget of 250 and 8 hours would be 260.

OpenStudy (gigi1103):

thanks!! can you help one of two more?

OpenStudy (texaschic101):

oh...I see now...my bad

OpenStudy (dangerousjesse):

Sure :) And no problem haha

OpenStudy (gigi1103):

k ill post it in a new question